OpenStudy (anonymous):

Find all solutions to the equation in the interval [0, 2π). cos 2x - cos x = 0

OpenStudy (tkhunny):

Have you any formulas for \(\cos(2x)\) in terms of \(\cos(x)\)?

OpenStudy (slaaibak):

change cos 2x to 2cos^2 x - 1

OpenStudy (anonymous):

Look at your trig identities

OpenStudy (anonymous):

cosx=1/secx

OpenStudy (anonymous):

i am looking. i understand that. i dont understand the solving part...

OpenStudy (tkhunny):

?? slaaibak already gave to you the one that will help.

OpenStudy (anonymous):

but why. i wanna understand.

OpenStudy (shubhamsrg):

cosa = cos b why don;t you generalize a simply ? In terms of b

OpenStudy (slaaibak):

because, when you only have a cos x 's, you can factorise it, which makes solving it easier

OpenStudy (anonymous):

because i'm not a genius. this is hard to teach yourself

OpenStudy (anonymous):

so now i have 2cos^(2)x-1-cosx = 0?

OpenStudy (tkhunny):

If it were this, \(2x^{2} - x - 1 = 0\), what would you do to solve it?

OpenStudy (anonymous):

-b (plus or minus) the square root of b^2-4(a)(c)/2a

OpenStudy (tkhunny):

Perfect. Do that with "cos(x)" instead.

OpenStudy (anonymous):

a=2cos^2x b=-1 c= nothing?

OpenStudy (tkhunny):

No. \(2\cos^{2}(x) - 1 - \cos(x) = 0\) or, in standard form, \(2\cos^{2}(x) - \cos(x) - 1 = 0\) a = 2 b = -1 c = -1 Now write the answer: \(\cos(x) = \dfrac{-b\pm\sqrt{b^{2}-4ac}}{2a}\) Do it EXACTLY like you would with \(2x^{2} - x - 1 = 0\) It is the same if you have \(2(Frog)^{2} - (Frog) - 1 = 0\). Write \(Frog = \dfrac{-b\pm\sqrt{b^{2}-4ac}}{2a}\) It is the same if you have \(2(y^{2})^{2} - (y^{2}) - 1 = 0\). Write \(y^{2} = \dfrac{-b\pm\sqrt{b^{2}-4ac}}{2a}\)

OpenStudy (anonymous):

\[\frac{ -cosx \pm \sqrt{cosx^2-4*2\cos^2x} }{ 2*2\cos^2x }\]

OpenStudy (tkhunny):

Why did you write that? Please go read that last response again. It starts out "cos(x) =" and cos(x) never appears again.

OpenStudy (anonymous):

cos(2x) - cos(x) = 0 ==> [2cos^2(x) - 1] - cos(x) = 0 ==> 2cos^2(x) - cos(x) - 1 = 0 ==> [2cos(x) + 1][cos(x) - 1] = 0 ==> cos(x) = -1/2 and cos(x) = 1.

OpenStudy (anonymous):

x = 0, 2π/3, 4π/3.

OpenStudy (tkhunny):

Absolutely, factoring is preferred to the Quadratic Formula.